OpenStudy (anonymous):

Compute the Derivative, if it exists: d/dx arcsin(x+1)^1/2

OpenStudy (anonymous):

www.themathpage.com/acalc/inverse-trig.htm tutorial.math.lamar.edu/Classes/CalcI/DiffInvTrigFcns.aspx

OpenStudy (anonymous):

im still getting a weird answer

OpenStudy (anonymous):

i get 1/(2-x)^(1/2)

OpenStudy (anonymous):

when the answer is 1/2((-x)^1/2)((1+x)^(1/2))

OpenStudy (anonymous):

\[1\div(2\sqrt{-x}\times \sqrt{1+x}\])

OpenStudy (anonymous):

That's almost what I'm getting. I used chain rule to get \[\large \frac{(1/2)(x+1)^{-1/2}}{1-(x+1)}\] Then simplified to get I used chain rule to get \[\large (-2x(x+1)^{1/2})^{-1}\] I might be missing something, though..

OpenStudy (anonymous):

and i use \[1\div \sqrt{1-(\sqrt{x+1){}}^2}\]

OpenStudy (anonymous):

you just plug it in like that, right?

OpenStudy (sirm3d):

don't forget the chain rule. you still have the derivative of \[\large (x+1)^{1/2}\] to account for in the final answer.

OpenStudy (anonymous):

ohh